# One-Versus-Others Attention: Scalable Multimodal Integration for Clinical Data

### v3

@@ Appendix B Simulation Dataset Details
Appendix C eICU Modality Descriptions and Detailed Pre-processing
The patient table contains patient demographic, admission, and discharge details. We use this table to determine mortality status. The diagnosis table contains active diagnoses given to each patient. We extract diagnosis features by one-hot encoding conditions specified in the ‘diagnosisstring’ column. The treatment table contains active treatments prescribed for each patient. We extract treatment features by one-hot encoding treatment types specified in the ‘treatmentstring’ column. The medication contains active medication orders for patients and when they were ordered. We extract medication features by one-hot encoding drugs specified in the ‘drugname’ column. The lab table contains lab type and corresponding results for various lab measurements collected for each patient. We extract lab features by summing lab measurements of commonly recorded lab types for each patient stay.
Appendix D TCGA Modality Descriptions and Detailed Pre-processing
CNV defines the varying number of repeats of genetic fragments found in a human genome. The number of repeats of specific genetic fragments influences gene expression levels and has been associated with the progression of different cancers [Samulin Erdem et al., 2017]. Any genomic regions missing CNV values or only having one unique value across all cases were removed. DNA methylation represents the amount of condensation of genetic regions due to the chemical alteration imposed by methyl groups. This condensation generally represses gene activity near the genetic region. Any genomic regions with missing values were removed. Clinical data includes information such as the patient’s diagnosis, demographics, laboratory tests, and family relationships. Categorical features were isolated and a coefficient of variation test was run to determine highly variable features. Features with a coefficient of variation higher than 70 were kept for analysis, along with the target variable. These features were converted into numerical format using one-hot-encoding. Gene expression data is collected through RNA-sequencing. Levels of gene expression are recorded by detecting the amounts of transcripts found for each gene. These levels can be used to determine the molecular mechanisms underlying cancer. Transcriptomic data was filtered to only include protein-coding genes and measured in fragments per kilobase of exon per million mapped fragments (FPKM). Imaging - TCGA collects pathology slide images of tissues sampled from the tumor. This modality provides visual information about the malignant region and can help with diagnosis and treatment planning. The image data was filtered only to include DX images, which result from a single X-Ray exposure, rotated to landscape view, then cropped to the median aspect ratio of 1.3565. We filtered for patients that had all five modalities, and we also only chose the patients that were still alive, to create a more balanced number of patients between cancer types (338 colon cancer patients, 329 kidney, 301 lung, 228 liver, and 226 stomach patients, after the filtering). The task we created is to classify each patient’s cancer type. For all modalities, features with missing values were dropped. For CNV, DNA Methylation, and gene expression data, feature reduction was performed using a random forest classifier, only on training data, ensuring the test was not seen by the random forest. Using the validation set, we determined the best number of estimators (out of 50, 100, 150).
Appendix E Hyperparameter Tuning
For each experiment, we used the validation accuracy to determine the best hyperparameters. We tuned the learning rate (0.01 - ), batch size (16, 32, 64, 128), epochs (200 epochs with early stopping if validation accuracy did not increase for 5 epochs), and number of attention heads for the OvO and pairwise cross-modal attention models (1, 2, 4, 8, 16). For the neural network encoders, we tuned the number of linear layers ranging from 1 to 4. Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4.
Appendix F Amazon Reviews Dataset Description and Detailed Pre-processing
The Amazon customer review dataset [Ni et al., 2019] aims to understand consumer opinions and experiences with products sold on Amazon. We sampled from the Electronics category, as it was one of the largest and most commonly used [Li et al., 2021, Haque et al., 2018, Pandey and Joshi, 2022]. We used three modalities from the dataset – review images, text, and the metadata (price, brand name, product category, etc.) associated with each review. The task is to classify the binary sentiment of each review - positive (4-5 stars) or negative (1-2 stars), with a total of around 20,000 samples. While sentiment classification with the Amazon reviews data is well-studied unimodally through text [Haque et al., 2018, Li et al., 2021, 2023, Pandey and Joshi, 2022], we offer a unique multimodal approach to the task. Existing multimodal approaches on Amazon reviews focus on review helpfulness [Han et al., 2022, Liu et al., 2021], or product similarity [Das et al., 2022] rather than sentiment classification. Furthermore, other papers use only text and images or only images and metadata (as text); an image-text-tabular multimodal task has yet to be explored. Intuitively, images and metadata can both be useful for review sentiment prediction - if the image highlights a broken item with the metadata showing a high price or trusted brand, consumer dissatisfaction may be implied.
Appendix G Compute Resources
@@
For each experiment, we use one NVIDIA GeForce RTX 3090 GPU. For the Hateful Memes task, single-modality models ran for roughly 40 minutes, and multi-modal models ran for roughly 55 minutes on average. For the Amazon reviews task, the single modality pre-trained models ran for roughly 50 minutes, the single modality neural network ran for a minute, and the multi-modal models ran for approximately an hour on average. For the TCGA task, single-modality models ran for 5 minutes, while multi-modal models ran for roughly 15 minutes on average. In the simulation dataset, the maximum modalities was 20 which took our model, OvO, roughly 2 minutes to run, while the cross-modal attention baseline took about 20 minutes to run on average.
Appendix H Significance Testing
@@
We use a t-test to determine if there is a significant difference the performance metrics (F1-score, accuracy, AUROC, AUPRC, MAUC, BCA) means between OvO attention and the next best-performing multimodal model. Our sample size is 10 from each group, as we initialized the models with 10 random seeds. For the MIMIC IV and CXR dataset, we compare against self-attention as it performed the second best after OvO. Using an = 0.01, we have evidence to reject the null hypothesis and conclude that there is a statistically significant difference in means between single-attention and OvO attention. The p-value for the AUROC scores is 0.00363 and the p-value for AUPRC is 0.000948. For the TCGA dataset, we do not have evidence to reject the null hypothesis and cannot say that the accuracy and F1-score means were different between OvO and cross-attention since the p-values were greater than = 0.01 (p-value of 0.04 for accuracy means, and p-value of 0.02 for F1-score means). This demonstrates that although cross-attention performed slightly better than OvO, it was not statistically significant. For the TADPOLE challenge, we compare against cross-attention as it performed the second best after OvO. We get a p-value for MAUC scores of and a p-value of for BCA. Thus, we demonstrate a statistically significant difference in accuracy and F1-score means between self-attention and OvO attention. Lastly, for the eICU dataset, we compare against cross-attention as it performed the second best after OvO. We get a p-value for AUROC scores of and a p-value of for AUPRC. Thus, we demonstrate a statistically significant difference in accuracy and F1-score means between self-attention and OvO attention.

### v4

@@ Appendix B Simulation Dataset Details
Appendix C Compute Resources
@@
For each experiment, we use one NVIDIA GeForce RTX 3090 GPU. For the MIMIC task, single-modality models ran for roughly 40 minutes, and multi-modal models ran for roughly 55 minutes on average. For the eICU, the single modality pre-trained models ran for roughly 50 minutes, the single modality neural network ran for a minute, and the multi-modal models ran for approximately an hour on average. For the TADPOLE task, single-modality models ran for 5 minutes, while multi-modal models ran for roughly 15 minutes on average. In the simulation dataset, the maximum modalities was 20 which took our model, OvO, roughly 2 minutes to run, while the cross-modal attention baseline took about 20 minutes to run on average.
[…]
Appendix D Significance Testing
@@
We use a t-test to determine if there is a significant difference the performance metrics (AUROC, AUPRC, MAUC, BCA) means between OvO attention and the next best-performing multimodal model. Our sample size is 10 from each group, as we initialized the models with 10 random seeds. For the MIMIC IV and CXR dataset, we compare against self-attention as it performed the second best after OvO. Using an = 0.01, we have evidence to reject the null hypothesis and conclude that there is a statistically significant difference in means between single-attention and OvO attention. The p-value for the AUROC scores is 0.00363 and the p-value for AUPRC is 0.000948. For the TADPOLE challenge, we compare against cross-attention as it performed the second best after OvO. We get a p-value for MAUC scores of and a p-value of for BCA. Thus, we demonstrate a statistically significant difference in MAUC and BCA means between self-attention and OvO attention. Lastly, for the eICU dataset, we compare against cross-attention as it performed the second best after OvO. We get a p-value for AUROC scores of and a p-value of for AUPRC. Thus, we demonstrate a statistically significant difference in AUROC and AUPRC means between self-attention and OvO attention.
[…]
Appendix E Limitations
This paper’s primary goal is to address one of the major challenges associated with multimodal datasets - computational resource demand and cost. While we demonstrate the scalability of the OvO attention mechanism and its efficiency in handling multiple modalities, we did not conduct experiments focused on interpretability, which is also a crucial aspect of multimodal learning in the biomedical space. The potential for OvO to reveal modality importance through the learninable parameter is promising, but further experiments are required to explore this interpretability in a meaningful way. Additionally, while we compared OvO to popular attention mechanisms like self-attention and cross-attention, there are other attention variants that could serve as future baselines, though they are not widely used in multimodal models. Lastly, the publicly available multimodal datasets often limit the diversity of modalities we can test, with many datasets primarily consisting of tabular data.